What is the average total cost of producing 40 units per day having cost(dollar per unit) of 1-8, and output (units per day) of 10-50. With the marginal cost on 40, average cost of 50, and the average total cost just in between these two.

1
Expert's answer
2022-01-10T09:55:15-0500

If Q = 50 and TC = 50×8 = 400, then ATC = TC/Q = 400/50 = 8.